OpenStudy (anonymous):

Find the derivative of sqrt(9-x)

OpenStudy (anonymous):

Remember that \[\huge √x = x^½ \ and \ 1/√x = x^-½ \] Also remember that \[\huge \frac {d(x^k)} {dx} = kx^{k-1} \]

OpenStudy (anonymous):

Get that

OpenStudy (anonymous):

-[sqrt(9-x)]^(-1/2) ???

OpenStudy (anonymous):

I understand that needs to be simplified, but is it essentially correct?

OpenStudy (no.identity05):